If the linear correlation between shoe size and height is 0.804 and Dave is 1.8 standard deviations above the mean in shoe size, what would you predict is Dave's height with respect to the mean height?
Answer #1

here as we know tht predicted z score zy=r*zx =0.804*1.8=1.45

hence we predict Dave's height to be 1.45 standard deviation above the mean height,